OpenStudy (anonymous):

\[z, A\in \mathbb C, b\in \mathbb R\] \[|z|^2+Re(Az)+b=0\text{ has a solution }\iff |A|^2\geq 4b\]

OpenStudy (jamesj):

This is a slightly obscure question. Is this a result you need as a lemma of a larger proof?

OpenStudy (anonymous):

this is the entire question

OpenStudy (jamesj):

and just to be clear, even though A is in caps, it is also just a complex number, yes?

OpenStudy (anonymous):

\[z=a+bi,A=e+fi\] \[a^2+b^2+ae-bf+b=0\] yes

OpenStudy (anonymous):

A in C say \[A=e+fi\]

OpenStudy (anonymous):

\[a^2+b^2\geq0\] \[ae-bf+d\leq0\] \[d\leq ae-bf\] stuck

OpenStudy (anonymous):

sorry i made a mistake in the question.

OpenStudy (jamesj):

Note that your last equation should be bf - ae \geq b.

OpenStudy (anonymous):

\[|z|^2+Re(Az)+d=0\]

OpenStudy (jamesj):

Ok and d is real.

OpenStudy (anonymous):

d real yes

OpenStudy (jamesj):

and i assume the inequality involves d, not b.

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

\[|A|^2\geq 4d\]

OpenStudy (jamesj):

I'll think about it during Doctor Who. :-)

OpenStudy (anonymous):

ty

OpenStudy (jamesj):

Ok, got it.

OpenStudy (anonymous):

@jamesj i sure don't love to see your answer...

OpenStudy (jamesj):

Write z = R.exp(beta.i) and A = S.exp(alpha.i) Then |z|^2 = R^2 and Re(zA) = RS.cos(alpha + beta) Thus |z|^2 + Re(zA) + d = 0 is R^2 + RS.cos(alpha + beta) + d = 0 and this quadratic has a solution if and only if the discriminent S^2 cos^2(alpha + beta) - 4d >= 0 i.e., S^2 >= 4d, as alpha and beta are arbitrary i.e., |A|^2 >= 4d

OpenStudy (anonymous):

thank you so much